Blog from February, 2022

Teardown day 5

Again fog and frost, but I digress....Clayton and I still are the primary ISFS crew.

Moved the reservation for the forklift from today to Wed.

Told Strawberry that we would return the 16' Penske tomorrow (after consolidating into the ISS Penske).

Applied ice melt around a few DCS structures blocking the trailer.

Drove up MH and removed (after a LOT of digging/pulling/prying) everything but 2 stakes holding the tower base plate.  Will have to return to this site (hopefully not driving).  Thanks Matt and Bill for helping!

Got lunch at the dairy.

With Dexter, pulled the DCS TT, still have one plywood pad frozen to the ground, but dug it out mostly.

Meantime, Liz, Bill, and Matt did the Penske consolidation.

And at the same time, Chris and Tony arrived after their Penske drive!

Remaining tasks:

  • DCS: pull out power cable, remove job boxes, darkhorses, transformers, plywood, remove fencing at 3 sites.  UU left a screw anchor at the site nearest the gate.  UU also left some anchors at the balloon site.
  • PRS: same as DCS, but also have to remove PRSG, drop the darkhorse, tower, and OTT stand.
  • MH: remove base plate
  • UP, MW, DC, CC, LC still need to be removed.  We had spent some time prepping CC cables and stakes.

Teardown day 4

Again, some fog in the west in the morning, truck temperature +3F, seemed to be linked to the Deer Creek reservoir.  Left a lot of fluffy frost on the DCS tower...

Spent the day at DCS with Clayton, Sebastian, and later Matt and Dexter.  Removed everything except the primary fence, the AC service, and one argumentative fence post and ground rod, preventing the trailer from being pulled.  Also some hardware (e.g. 2 darkhorses and the Rohn sections) is still there.

Extracted all soils with a few nicks, but presumed working.  The heat flux plate was off of horizontal by perhaps 5 degrees and the TP01 foil was a bit wrinkled.

At the very end of the day, ran out to CC to soak the lock with Kroil.  It now works with the combination we expected.

Fog 26 Feb

Fog this morning along the entire west side of the valley.  SH, PC, and PRS are now down, along with all(?) of ISS, but perhaps there are some interesting data?

Teardown day 3

Fog this morning, but the program is over....  Truck thermometer said -1F when we drove out.

Started the day removing the TT trailer from PRS, along with remaining sensors.  A bit of an effort:

  • T post and ground rod required the Hi-Lift.  We rigged up ways to attach to them.
  • Even after cranking the outrigger legs up, there was still a lot of tension on one.  Used the Hi-Lift again to remove the foot from this outrigger
  • Pushing the outriggers in, only one went all the way.  The other 3 got stuck with about a foot remaining and wouldn't extend back out either.  We suspect something is frozen in there.  During the drive out on washboard, one of the 3 became loose and was able to be reseated.  We'll try to reseat the others once the weather warms.

Next removed all of SP.  Not TOO bad, though a stake and ground rod required the Hi-Lift again.  Recovered all 3 power cables and all 4 soil sensors.

Spent a bit more time at the base and Penske organizing.

Finally, at CC, dug out all of the cables, tops of T stakes, and removed snow cover from the soil plot, with some help from Liz, Matt, and Dexter.  A lot of ice in spots – especially by the job boxes – and none of the 8(?) T stakes that I tried budged.  My intent was to prepare the site to have a quick removal with the forklift on Monday.  Now, I'm not so sure.  Oh, and we couldn't get the gate lock to open with the combo I had photographed on the marker board.  I need to see if I can find the original email from the manager.

Teardown day 1

Last IOP went until 8AM.  They got a good fog case and were very happy.

Removed SH

  • broke the foil from the second "extra" TP01 – it just snapped off in the cold
  • nicked the cable of the Qsoil, which will need to be repaired
  • damaged 3? of the power cables while extracting them from the ice

Removed PC

  • also broke the foil on the TP01, even though I had mostly freed it.  Grrrr.
  • also nicked one power cable.  Cut the other 2 power cables where Dan had buried them under snowmobile tracks
  • Tsoil appeared to be deeper – the top about 1.5cm below the soil, though it is hard to distinguish dirt and frozen biomass
  • Gaoil came out tilted by about 30 degrees off vertical.  Might have been in the extraction, or might have been as the soil thawed/refroze during the project.

A long day just doing this, but progress....

Teardown day 2

Clayton, Sebastian, and I almost completely removed instrumentation from PRS.  The only sensors standing are the NR01, CS125, and a ceilometer, mostly because we didn't have shipping boxes. Almost all cables (except the main power) have been removed.  The TT is down and nearly ready to be pulled out tomorrow. 

Was able to remove all soil sensors, in part because the ground never really healed after the install, so I was able to break open several of my "clods".

A really good day!

Status 23 Feb

Last (full) day of ops!  IOP9 will officially end at 0800 tomorrow.  Clayton is now here (and Liz, Lou, and today Matt for ISS).

Snow off and on all day, another few inches of accumulation.  Stopped at 1630, but now (2030) flurries again.

Spent day orienting Clayton and tear-down prep.  Picked up Penske and started filling with spare stuff going to Boulder and empty boxes for tearing down satellite sites tomorrow.  Grabbed a few more tools for the pickup.  Reserved an all-terrain forklift for Monday to pull out CC.

At the end of the day, reinstalled a bunch of thermocouples at DCS.  Lowered the tower to replace its thermocouples, but attempts to tilt blew the breaker on the generator.  We presume that the hydraulic pump is just too cold.  We raised the tower again, with no change to the instruments (of course).  Tower down from about 1720-1820.  I hope the thermocouples last!

BTW, noticed that the weather service radar (based on the one at SLC) is poor at detecting light snow.  The radar image showed that snow had passed Heber City by 2:30 today, when it was clearly snowing out my window at the Best Western.

Status 22 Feb

Snow stopped and clouds broke up about 8AM.  A nice, sunny day for most of the morning.  Terry left about 15min before this happened, so never saw the mountains!

Bought tire chains and checked out snowmachine rental.

DCS: broke up a bunch of ice, freeing the main power cable.

Base: organized a bit, then a bit more!

Helped Bill with the 4pm sounding, since there was some wind and there were some gas issues.

Status 21 Feb

The forecasted snow did come, starting about 5AM and continuing off and on through the day.  Accumulation about 2".

Spent the day touring with Terry and later Erik and Sebastian:

  • ISS1: Bill showed us his site
  • DCS
  • Balloon site
    • brushed off snow from Sebastian's Apogee
    • shifted tethersonde back to its donut and brushed off some snow
  • SH (then ate lunch at the SH grill)
    • sonic is tilted down – I'm guessing that my shim washer slipped out.  We should check the time history of tilts
  • ISFS base, where we met up with E&S, who had never seen some of the satellite sites...
  • DC
    • noticed a drip of water off the top EC150 window
  • CC
  • LC
    • cleaned bird doo from
  • SP
    • entire sonic boom also leaning.  In part because of the T-stake being broken on the opposite guy, so there isn't as much tension.  Again, look at tilts
  • MW
  • drove past the base of MH
  • Back to ISS1, where Bill was cleaning off snow
Status 20 Feb

Jacquie and John departed, Bill is now here.  Thanks, so much, J&J!

Weather was clear, reasonably warm, but expecting snow tomorrow.  Ahead of the storm was a gust front at about 11:30 where speeds at DCS had a step change from 2 to 9 m/s.

No data issues starting out – the replacement CSAT at LC worked all night.  Nevertheless, I did the "ultimate" EOL/CFACT, by visiting EVERY EOL location today.  Since snow is predicted, probably through the end of the project, Sebastian requested that I pull at least the thermocouples at the satellite sites...

  • Started at the ISFS base, where I bonked my head
  • DCS: attached a license plate to the trailer
  • Back to base to participate in the CFACT zoom – no IOPs for at least the next 2 days
  • SH: remove thermocouple
  • ISS trailer: meet Bill and pick up ladder
  • Lunch
  • DC: remove thermocouple
  • CC: remove thermocouple (snow gets deeper – about 1' – out in the field
  • LC: remove thermocouple (used stepstool beyond specifications)
  • UP: remove thermocouple (field is driveable)
  • PC: remove thermocouple
  • MH: remove thermocouple
  • PRS: attach license
  • MW: remove thermocouple (used tall ladder)
  • SP: remove thermocouple (one T-stake is missing its T)
  • Balloon trailer: saw Bill just after the sonde launch
  • Back to base to tidy up.

Terry arrived in time for dessert!

Ops 2/19

Weather:  Mostly sunny and calm. Highs in the mid 40's. 

IRGA at the supersites - Noticed this morning that there was a long period of bad IRGA data at the super sites at certain heights. Up to 12 hours of bad data at some heights - longer than I’ve seen in the time series. In particular, heights

  • irga.1m.prs
  • irga.3m.prs
  • irga.2m.dcs
  • irga.7m.dcs

Zoom in on the past day

Past week

QC table screenshot 


  • Steve double checked that the heaters were set properly.
  • We visited the DCS site to check the sonic/irga at 2m but there was nothing unusual to note.
  • Steve noted that odd HRXL data may be due to interference in the field-of-view with the boom that holds the sonic/irga. I noted that in the DM tracker. We should do a closer examination of these time series during post-project data QC.

  • While at DCS we met up with Eric and Sebastian is it looks like they successfully captured early morning fog during IOP8. They had just finished operating the tethered sonde. John and Steve helped secure the balloon


  • Steve visited the LC site to swap the bad sonic with Sebastian's (See Steve's Blog post).  I made a note in the DM tracker. 


Mid-afternoon we did a second check of the data and sensors at ISS1 then drove to the radiosonde site to launch the 415p radiosonde. We had a pleasant surprise when the PI Zhaoxia Pu from U of Utah stopped by. She assisted filling the balloon and launching the afternoon radiosonde - see John's ISS blog post.  

Today is the last day John and I are supporting CFACT operations in the field. Sunday, Steve and Bill take over (Bill arrives tonight). It has been eventful! 

Signing off. Time for supper!

Swapped sonic at LC

Because the lc sonic has been dying during the night the last weeks(?), and a swap of the EC100 didn't fix it, I've just swapped in Sebastian's CSAT3A (old style).  I just placed a bolt through our cassette mount, that is a bit sketchy, but should hold unless we get strong winds.  Also, I didn't figure out an easy way to move the EC150 (as I had done for SH), so the sonic centerline is now higher than the EC150 midpoint by perhaps 5cm.

I started with 3 greens (since it was daytime) and ended with 3 greens, so I guess I did no harm....

Ops 2/18

Weather: Mostly sunny, calm. Highs in the mid 40’s. 

Data checks today revealed no glaring issues as of this blog. 

IOP08 today from noon Friday - noon Saturday

  • PI’s called another IOP for today on the heels of IOP07. John and I launched the 2p and 4p radiosondes. 
  • Otherwise a rather quiet day. 

Steve arrived today! 

Ops 2/17

Weather: Mostly sunny. Calm to breezy winds from the South. Highs in the upper 30’s. 

IOP07 was called today from noon Thursday - noon Friday

We first did our typical morning data check walk-through at ISS1 and the base trailer. Then off to the supersites to replace thermocouples and clean radiation sensors. 


Replaced the thermocouple at t.7m. The towers were down from 1000 - 1111. John was the tallest and cleaned all the towers radiometers with the aid of a ladder, plus the ground-level one’s to take advantage of having the methanol/kimwipes on hand. 

Note that the generator fuel is down to a 1/4 of a tank:


Replaced the thermocouples at t.3m and t.32m and cleaned all the radiation sensors. Towers were down 1130 - 1235. We had a large impromptu gathering of curious cows in the meantime. Initiating the car alarm worked wonders clearing the audience.

Note the generator fuel is half full. 

Sebastian mentioned he visited PRS soon after us to install the ground-level thermocouples and had to re-leveled the radiometer again.  It keeps drooping.

We launched the radiosondes at 210p and 415p (see John’s ISS blog for details).

Finally, Sebastian found a license missing from the front of the Silverado truck. The cows probably knocked it off. I knew they were up to no good! It's now with the truck.

Ops 2/16

Weather: Mostly cloudy and overcast. Definitely a chillier and moist day. Light winds mostly from the north. High's in the upper 30's.

A quiet data day for both ISFS and ISS. 

While launching the 415p radiosonde Sebastian came by to ask about the remaining radiosonde/helium inventory (see John’s ISS blog post for details).  

Sebastian mentioned he has had to re-level the radiometers mounted on the dark horse which have started drooping. This must be a recent event because while John and I last visited those sites we didn’t notice any sagging. Today he re-leveled the one at DCS ~350-410p. He also cleaned the radiometers with methanol and kimwipes at 0.5m and 2m at DCS and made plans to do the same at PRS by the end of today. 

We discussed the thermocouples at the supersites and I noted that a few of the thermocouples were broken at the DCS and PRS sites since this morning:

  • t.7m@DCS 
  • t.3m (rohn), t.32m (tt) @PRS

Plan for tomorrow

There is a tentative IOP7 planned for noon tomorrow (Thursday) to noon Friday. So we plan to lower/raise towers at the DCS and PRS sites in the AM to replace the broken thermocouples. We're bringing the ladder with us to replace the t.3m on the rohn at PRS. In addition, Sebastian suggested that we clean the radiometers at both towers - he has left us with kimwipes and a bottle of methanol. We will aim to do that as well.